OpenStudy (lgbasallote):

Prove that if x is irrational then 1/x is irrational

OpenStudy (lgbasallote):

i suppose first step is to assume x is rational and try to prove by contradiction...

OpenStudy (lgbasallote):

so x = a/b where a/b is in simplest terms

OpenStudy (lgbasallote):

nevermind....this is wrong....

hartnn (hartnn):

i would go by contadiction, let us assume x is irrational, then 1/x is rational.(we'll prove this incorrect) if 1/x is rational, what it can be written as?

OpenStudy (lgbasallote):

i assume b/a where b/a is in simplest form

hartnn (hartnn):

same thing in different words, 1/x = b/a where a,b are integers then what is x from here?

OpenStudy (lgbasallote):

a/b

OpenStudy (lgbasallote):

oh...

OpenStudy (lgbasallote):

i see

hartnn (hartnn):

since, a and b are integers, a/b = x is rational

OpenStudy (lgbasallote):

is it possible to do contradiction using x = a/b?

hartnn (hartnn):

then u will prove that if x is rational, 1/x is also rational

OpenStudy (lgbasallote):

shouldn't it be 1/x is irrational?

OpenStudy (lgbasallote):

oh...i see

hartnn (hartnn):

if x is rational, 1/x is also rational